In an experiment it was found that when a sonometer in its fundamental mode of vibration and a tuning fork gave 5 beats when length of wire is 1.05 metres or 1 metre. The velocity of transverse waves in sonometer wire when its length is 1 metre

Grade:12

1 Answers

Arun
5 years ago
Dear student
 
As we know
 
n + 5/n- 5 = L2/L1 
 
n+5 /n-5 = 1.05/1
 
0.05 n = 10.25
 
n = 10.25/0.05 = 205 Hz
 
Hope it helps
 
n = 4.75/0.05 =